1. A hot air balloon has a volume of 2800 m3 at 990C. What is the volume if the air cools to 800C.

Given:     V1 = 2800m3

T1=99°C+273K = 372 K        T2 =80°C +273K =353K

Unknown: V2

Formula:

Solution:

Final answer:



1
Expert's answer
2022-05-03T08:58:04-0400

Given:

V1 = 2800 m3

T1 = 99°C + 273.15 = 372.15 K

P1 = P2 = constant

V2 = ???

T2 = 80°C + 273.15 = 353.15 K


Formula: V1 / T1 = V2 / T2


Solution:

Since the pressure and amount of gas remain unchanged, Charles's law can be used.

Charles's law can be expressed as: V1 / T1 = V2 / T2

V1T2 = V2T1

To find the volume, solve the equation for V2:

V2 = V1T2 / T1

V2 = (2800 m3 × 353.15 K) / (372.15 K ) = 2657 m3

V2 = 2657 m3


Answer: The volume is 2657 m3